R3al:

Simplify 5/8 + 3/4 + 2/5

carmelle:

First we need to find the common denominator for 8 4 and 5

R3al:

ok

carmelle:

do u have an idea of what it can be?

R3al:

so do we 8x4x5 ?

supie:

\[5/8+3/4+2/5=\]\[11/8+2/5=?\]

R3al:

I still need help, So do we add the top numbers?

carmelle:

no the bottom ones

carmelle:

he simplified it

R3al:

so you add the bottom ones?

carmelle:

u need a common denominator tho

R3al:

ok

carmelle:

yea.. I'd say 40 is a common denominator...

R3al:

ok but how did you get that?

carmelle:

8x5

R3al:

oh, ok.

carmelle:

we are doing it

R3al:

Ok,since we have the common denominator, what do we do next?

carmelle:

11/8 = 55/40, since we multipiled the denominator by 5, we have to do the same with the numerator

R3al:

ok

carmelle:

2/5 = /40 Do u know what the numerator should be?

R3al:

10?

carmelle:

no.. try again... multiply by 8

R3al:

so 10 x 8 = 80

carmelle:

no.. multiply 2 by 8

carmelle:

and what do u get

R3al:

16

carmelle:

correct

carmelle:

now we can add..

R3al:

ok

carmelle:

55/40+16/40 is equal to..

R3al:

71/40?

carmelle:

yup that's correct

R3al:

Really, dang!

carmelle:

do u need it as a mixed fraction.. or no?

R3al:

yes

carmelle:

do u know how to get it as a mixed fraction?

R3al:

Divide the numerator by the denominator?

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @R3al Divide the numerator by the denominator? \(\color{#0cbb34}{\text{End of Quote}}\) we are subtracting

carmelle:

71 - 40

R3al:

31

carmelle:

yea so it would be 1 31/40

carmelle:

\[1\frac{ 31 }{ 40 }\]

R3al:

Oh, So that's the Answer?

SageG:

Say what now? Subtracting?

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @R3al Oh, So that's the Answer? \(\color{#0cbb34}{\text{End of Quote}}\) yup 71/40 or 1 31/40

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @SageG Say what now? Subtracting? \(\color{#0cbb34}{\text{End of Quote}}\) no we are done lol

SageG:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle \(\color{#0cbb34}{\text{Originally Posted by}}\) @SageG Say what now? Subtracting? \(\color{#0cbb34}{\text{End of Quote}}\) no we are done lol \(\color{#0cbb34}{\text{End of Quote}}\) No i mean to get a mixed fraction

R3al:

Thank you SO MUCH @carmelle !!!

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @SageG \(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle \(\color{#0cbb34}{\text{Originally Posted by}}\) @SageG Say what now? Subtracting? \(\color{#0cbb34}{\text{End of Quote}}\) no we are done lol \(\color{#0cbb34}{\text{End of Quote}}\) No i mean to get a mixed fraction \(\color{#0cbb34}{\text{End of Quote}}\) yea... that's what I was taught

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @R3al Thank you SO MUCH @carmelle !!! \(\color{#0cbb34}{\text{End of Quote}}\) u r very welcome :)

SageG:

I see, i had bo idea thats how you get it. I use to divide and or just wing it tbh but avoid it usually

SageG:

No*

SageG:

Btw nice work

carmelle:

ty (: